OpenStudy (anonymous):

if a yard has an area of 96 sq ft .. the width of the yard is 4 ft less than the length what is the perimeter in ft of the yard how can i solve this ?my head is in knots

OpenStudy (ivettef365):

4 ft less than the length so it is L - 4

OpenStudy (ivettef365):

Formula is P = 2L + 2W W = (L-4) so 96 = 2L + 2(L-4) solve for L

OpenStudy (anonymous):



OpenStudy (anonymous):

If pictures help?

OpenStudy (anonymous):

i understand that but how can i use 4 an d96 to solve

OpenStudy (anonymous):

Like she said before you make an equation equal to 96 using the variable L nd -4.

OpenStudy (anonymous):

You agree with there being two sides equalling to L and two sides equaling to L-4 right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So make an equation out of that....

OpenStudy (anonymous):

2L+2(L-4)=96

OpenStudy (anonymous):

ok il try solving thx

OpenStudy (anonymous):

But remember your not finished after that! you have to add up all the sides to get the perimeter.